F(x)=x^3-x^2-8x+12

Write the values of x for which f'(x)>0


1
Expert's answer
2021-05-31T18:25:53-0400

"f(x)=x^3-x^2-8x+12."

"f'(x)=3x^2-2x-8=(3x+4)(x-2)."

"f'(x)>0" when "-\infin<x<-\frac{4}{3}" or "2<x<\infin."